STATISTICS. CONFIDENCE INTERVALS.

Let \left ( X_{1},X_{2},...,X_{n} \right ) be a simple randon sample of a population with distribution  N\left ( \theta ,\sigma ^{2}= 100 \right ).

Construct a credible region with probability 0.95 for the mean \theta , if it is assumed that initial distribution for \theta is N\left ( \mu _{0}= 100,\sigma _{0}^{2}= 225 \right ).

Thank you for your explanations.
Answer #1

here, mean = 100
and std. dev. = sqrt(225) = 15

z-value for 95% CI is 1.96 (here z-value indicates the area in the left/right tail is 0.025)

CI = (mean - z*sigma, mean + z*sigma)
= (100 - 1.96*15 , 100 + 1.96*15)
= (70.6 , 129.4)